[diffusion] Diffusion Morning Report 03/01/13

Ted Berg tberg at stanford.edu
Fri Mar 1 06:13:56 PST 2013


Good Morning All,
*1**. New issues**:* None overnight
* 2. Ongoing issues:*Tylan Poly rewriting recipes,** Tylannitride  
tested again waiting further instructions after alarms, *O2 in Silane on 
?EPi, * EPI Chamber A *pump out for repair,**Tylan 3 thin,**Tylan 2 poor 
uniformity*, Thermco 1 thickness issues?
*3. Resolved:**Epi Chamber B running Fine

*
					
	Date: 	3/1/13 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*needs test* 	** 	** 	**
Tylan 3 	*yellow* 	*thin?* 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*yellow* 	*temp issue-front zone low* 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*yellow* 	*needs test* 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*waiting for users* 	*3.4/30* 	** 	**
Tylan 9-Poly 	*down* 	*rewriting recipes-testing revised recipes* 
*23.7/30* 	** 	**
Tylan 10-Nitride 	*yellow* 	*for anneals* 	*28.5/30* 	** 	**
Tylan 11-LTO 	*yellow* 	** 	*4.3/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*yellow* 	*thin?* 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*up* 	** 	** 	** 	**
Thermco4 	*up* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*down* 	*pump out for repair-hopefully back today* 
** 	** 	*1*
AMAT Epi Cham. B 	*yellow* 	*testing* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	***changeover working 
well*** 	** 	**

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20130301/a6fe8e7b/attachment-0001.html>


More information about the diffusion mailing list